# When Geoscience Meets Foundation Models: Towards General Geoscience Artificial Intelligence System

### v3

@@ III-A2 Adaptation
RingMo-SAM [52] Remote sensing images Generative Model fine-tuning Remote sensing image segmentation Effective across various datasets
SAMRS [68], RSPrompter [36]
Remote sensing images
Prompt engineering
Remote sensing image segmentation
Segment Anything Model (SAM)
@@
Ma et al. [69] Remote sensing images Generative Model fine-tuning Remote sensing image segmentation Segment Anything Model (SAM)
Geo-SAM [70]
Remote sensing images
Model fine-tuning
Infrastructure segmentation
Segment Anything Model (SAM)
@@
OFA-Net [71] Remote sensing images - Model fine-tuning Infrastructure segmentation Segment Anything Model (SAM)
CSP [72]
Remote sensing images
Contrastive
Model fine-tuning
Satellite image classification
Satellite
@@
Prithvi [16] Satellite imagery Generative Model fine-tuning Earth observation tasks open source
ClimaX [73], W-MAE [74]
Grid-based meteorological data
Generative
Model fine-tuning
Climate change prediction
Climate and weather modeling
@@
Multi-modal Remote Sensing Imagery dataset
Contrastive
Model fine-tuning
Earth image analysis
State-of-the-art performance
LVLMs
RSVQA[76]
Images and text
Generative
Model fine-tuning
Satellite image analysis
Satellite image analysis
@@
Chappuis et al. [78]
Exploration of language biases in RSVQA
@@
EarthGPT [80]
Image-text pairs
Instruction tuning
Universal RS image comprehension
Open-set reasoning tasks.
@@
SkyeyeGPT [81] RS instruction-following dataset - Instruction tuning Image-level and region-level tasks Comparable results with GPT-4V
Mall et al. [82]
Multimodal geospatial data
Generative
Model fine-tuning
Geospatial data analysis
Satellite images
@@
RS-CLIP [83] Image–text datasets Contrastive Model fine-tuning Remote sensing scene classification Zero-shot remote sensing scene classification
Zhang et al. [84]
Images and text
@@
Evaluation of LVLMs in Earth observation tasks

### v4

@@ IV-B3 Applications and Discussions
RingMo [18], Wang et al. [19], Cha et al. [214] Remote sensing images Generative Fine-tuning Remote sensing image analysis -
RingMo-Sense [215] Remote sensing images Generative Fine-tuning Spatiotemporal prediction Video Swin Transformer
RingMo-Lite [216] Remote sensing images Generative - RS image interpretation Optimized for edge devices
@@
RingMo-SAM [34] Remote sensing images Generative Fine-tuning Remote sensing image segmentation Effective across various datasets
SAMRS [217], RSPrompter [218] Remote sensing images - Prompt tuning Remote sensing image segmentation SAM
@@
Ma et al. [219] Remote sensing images Generative Fine-tuning Remote sensing image segmentation Segment Anything Model (SAM)
Geo-SAM [220] Remote sensing images - Fine-tuning Infrastructure segmentation Segment Anything Model (SAM)
OFA-Net [230] Remote sensing images - Fine-tuning Infrastructure segmentation Segment Anything Model (SAM)
CSP [221] Remote sensing images Contrastive Fine-tuning Satellite image classification Satellite
@@
Prithvi [20] Satellite imagery Generative Fine-tuning Earth observation tasks open source
ClimaX [222], W-MAE [223] Grid-based meteorological data Generative Fine-tuning Climate change prediction Climate and weather modeling
@@
SkySense [224] Multi-modal Remote Sensing Imagery dataset Contrastive Fine-tuning Earth image analysis State-of-the-art performance
Shankar et al. [231, 232] Sentinel-1 and Sentinel-2 imagery Contrastive Fine-tuning glaciological image analysis -
Julka et al. [227] HiRISE Images Hybrid Remapping Tuning Planetary image analysis SAM
Osco et al. [233] Remote sensing images Hybrid Fine-tuning Satellite imagery analysis SAM
Sun et al. [226] Satellite imagery Hybrid Fine-tuning Parcel segmentation SAM
RSbuilding [225] Satellite imagery Hybrid - Building extraction and change detection -
VSGNet [228] Multi-modal RS images - Fine-tuning RS semantic segmentation -
Satvit [48] Satellite-derived RS images Contrastive Fine-tuning RS images analysis ViT
SpectralGPT [49] Spectral remote sensing images Generative Fine-tuning captioning and retrieving remote sensing images 600 million parameters
Rsmamba [26] remote sensing images - - remote sensing image classification -
Ren et al. [234] overhead imagery - - various overhead imagery tasks SAM evaluation
DF4LCZ [235] Google imagery & multispectral imagery - - local climate zones classification -
Gallagher et al. [236] Multispectral imagery - Prompt tuning multispectral object detection SAM
S2MAE [229] Spectral RS images - - - A million training images
Feng et al. [237] Cross-modal data - - Multiple downstream tasks
IV-C Large Vision-Language Models
Figure 15: CLIP [142].
IV-C1 Basic Large Vision-Language Models
In the current landscape, FMs have evolved to encompass more than just deep models; they now focus on utilizing vast amounts of data and computational power to tackle diverse problems. The goal is to use a unified model capable of addressing multiple modalities and tasks. Consequently, there is a growing emphasis on training vision-language models [238].
CLIP [142] is a prominent example that leverages a large-scale collection of image-text pairs from the Internet for contrastive learning, enabling the creation of a unified representation of multimodal data. Inspired by CLIP, various multimodal image FMs have been proposed, differing in their model structures [239, 240], feature representations [241, 242], multimodal feature fusion approaches [243], feature alignment loss functions [244], pretraining methods [245], and more. These advancements have significantly improved the performance of multimodal FMs.
In addition, DALL-E [136] combines the diffusion model with multimodal FMs to generate images from text. SAM [213] introduces a promptable model with training strategies that enable the segmentation of objects using text, points, and lines as prompts. GPT-4 [36] exhibits superhuman capabilities on various professional and academic datasets. mPLUG-2 [246] introduces a multimodule composition network, including text, image, and video. Metatransformer [247] proposed a unified framework performing learning across 12 modalities with unpaired data (e.g., natural language, 2-D images, 3-D point clouds, audio, video, time series, and tabular data).
IV-C2 Advanced Large Vision-Language Models for Geoscience
LVLMs have emerged as powerful tools in remote sensing and geoscience, particularly for visual question answering and image understanding tasks. Chappuis et al. [248] pioneered the use of Remote Sensing Visual Question Answering (RSVQA), leveraging advancements in natural language processing and computer vision to analyze satellite images. This work was further developed by the work [249], who introduced LiT-4-RSVQA, a lightweight transformer-based architecture for efficient and accurate visual question answering in remote sensing. However, the authors in [250] later identified the need to address language biases in RSVQA, revealing more severe biases in remote sensing compared to standard VQA due to dataset specifics.
In the realm of remote sensing image understanding, several notable developments have occurred. Kuckreja et al. [251] introduced GeoChat, the first remote sensing VLM with multitask conversational capabilities. The authors in [46] and the work [252] presented EarthGPT and SkyeyeGPT respectively, both LVLMs tailored for Earth observation and remote sensing tasks. To address the lack of comprehensive datasets, Hu et al. [253] developed a high-quality remote sensing image captioning dataset (RSICap) and a benchmark evaluation dataset (RSIEval).
Evaluation and benchmarking of LVLMs in remote sensing have been crucial for understanding their capabilities and limitations. The authors in [254] conducted a comprehensive benchmark focusing on scene understanding, localization, counting, and change detection tasks, revealing that while these models excel in tasks like location understanding and captioning, their spatial reasoning limitations affect performance in object localization and counting. Li et al. [255] introduced VRSBench, a comprehensive vision-language benchmark for remote sensing image understanding, featuring a large dataset of verified captions, object references, and question-answer pairs.
Recent research has focused on developing more advanced LVLM architectures and techniques specifically tailored for remote sensing applications. Liu et al. [256] presented RemoteCLIP, a vision-language foundation model that leverages data scaling and incorporates UAV imagery. The work [257] developed CLIP-RS, a cross-modal remote sensing image retrieval platform. The authors in [258] introduced H2RSVLM, a model trained on a large dataset for enhancing self-awareness in remote sensing image understanding tasks.
Further advancements include the work [259], who explored the use of CLIP/Open-CLIP variants for zero-shot classification in remote sensing, and He et al. [260], who proposed a foundation model adaptation (FMA) framework for multimodal remote sensing data classification. The authors in [237] developed a remote sensing cross-modal foundation model that uses multi-domain representation and cross-domain fusion to extract generalized features from massive multi-modal data.
IV-C3 Applications and Disscussion
LVLMs have found diverse applications in remote sensing, including visual question answering [248], scene classification [261], object detection, change detection, image captioning, and multi-modal data fusion [237]. These applications have the potential to revolutionize fields such as urban planning, environmental monitoring, and disaster response. However, the advancement of LVLMs in remote sensing also presents challenges. While models like GeoChat [251] and lightweight architectures like LiT-4-RSVQA [249] improve accessibility and efficiency, issues of bias [250], interpretability, and ethical use remain critical concerns. The development of high-quality datasets [253] and benchmarks [255] is crucial for progress, as is the integration of domain-specific knowledge. As LVLMs continue to evolve, balancing their immense potential with careful consideration of fairness, privacy, and security will be essential to their successful application in Earth observation and geospatial analysis.
The rapid advancement of LVLMs in geoscience and remote sensing presents both opportunities and challenges for the field. While models such as GeoChat [251] and RemoteCLIP [256] demonstrate the potential to democratize access to Earth observation data and improve performance across various tasks, they also raise important concerns. The work by Chappuis et al. [250] on language biases in RSVQA highlights the need for careful consideration of fairness and bias mitigation, especially given the global nature of satellite imagery. As these models become more complex and widely used, ensuring interpretability and trustworthiness becomes crucial, particularly for applications in critical areas such as disaster response or environmental monitoring. The development of high-quality datasets and benchmarks, as seen in the work of Hu et al. [253] and Li et al. [255], is essential for advancing the field, but creating large-scale, diverse, and accurately labeled datasets remains a challenge. Furthermore, as LVLMs become more powerful in analyzing high-resolution satellite imagery, addressing privacy and security concerns becomes increasingly important. Future research should focus on integrating domain-specific knowledge, improving model generalization, and developing ethical guidelines for the responsible use of these technologies in Earth observation and geospatial analysis.
TABLE IV: Summary of Advanced Large Vision-Language Models for Geoscience
Ref. Data Type Pre-training Adaptation Applications Key Features
RSVQA[248] Images and text Generative Model fine-tuning Satellite image analysis Satellite image analysis
@@ IV-C3 Applications and Disscussion
Chappuis et al. [250] - - - - Exploration of language biases in RSVQA
@@
EarthGPT [46] Image-text pairs - Instruction tuning Universal RS image comprehension Open-set reasoning tasks.
@@
SkyeyeGPT [252, 262] RS instruction-following dataset - Instruction tuning Image-level and region-level tasks Comparable results with GPT-4V
Mall et al. [263] Multimodal geospatial data Generative Fine-tuning Geospatial data analysis Satellite images
@@
RS-CLIP [261] Image–text datasets Contrastive Fine-tuning Remote sensing scene classification Zero-shot remote sensing scene classification
@@
Zhang et al. [254] Images and text - - - Evaluation of LVLMs in Earth observation tasks